OpenStudy (anonymous):

One side of a rectangle is 3 inches shorter than the other side, and the perimeter is 54 inches. What are the dimensions of the rectangle? Which of the following equations could be used to solve the problem? 2x + 2(x - 3) = 54 2x + 2x - 3 = 54 2x - 3 = 54

OpenStudy (anonymous):

2x-3=54

OpenStudy (anonymous):

2x + 2(x - 3) = 54 x = one side x - 3 = other side 2x + 2(x - 3) = perimeter

OpenStudy (anonymous):

im2 is wrong

OpenStudy (anonymous):

....omg u aint gotta make it a big deal

OpenStudy (anonymous):

<3 lul @lm2

OpenStudy (anonymous):



OpenStudy (anonymous):